4.We draw 7 cards from a well shuffled deck of 52 cards.


a) What is the probability that the 7 cards include exactly 3 aces?

b) What is the probability that the 7 cards include exactly 2 queens?

c) What is the probability that the 7 cards include exactly 3 aces,or 2
queens,or both?

Answer by stanbon(75887)

4.We draw 7 cards from a well shuffled deck of 52 cards.

a) What is the probability that the 7 cards include exactly 3 aces?
Pick 3 aces: 4C3 = 4C1 = 4 ways
---
Pick 5 other cards: 48C5
---
Probability = (4*48C5)/52C7 = 0.0512
-----------------------------------------
b) What is the probability that the 7 cards include exactly 2 queens?
Prob = (4C2*48C5)/52C7 = 0.0548
--------------------------
c) What is the probability that the 7 cards include exactly 3 aces,or 2
queens,or both?
Prob = 0.0512+0.0548 + [(4*6*44C2)]52C7 = 0.1062
==================================================
Cheers,
Stan H.
==============